# Fast Dynamic 3D Object Generation from a Single-view Video

### v1

@@ 4.3 Quantitative evaluation
In Tab. 1, we compare Consistent4D and our method based on average generation time, rendering time, and CLIP-similarity scores [31] for input image sequences and rendered images. Measurements used the image sequences described in Sec. 4.1. Consistent4D’s CLIP-similarity evaluation is limited to RGB images from 2-3 selected views in their paper and project page, while our evaluation includes 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps, making our CLIP-similarity score more robust.
We draw several observation: (1) Our method achieves superior image quality. (2) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement. (3) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization in fewer iterations, and (2) each iteration in our method requiring less time compared to Consistent4D. This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
@@
Input Generated images Geometry
Figure 5: Given only two input frames, our method is able to generate point clouds and temporal continuous images from novel views.
@@ 4.4 Sparse input evaluation
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In contrast, Consistent4D fails to operate effectively under such conditions. In such cases, we set $w=0.25$ in Eq. 2. As illustrated in Fig. 5, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
@@ 4.5 Ablation studies
Generated images Reconstructed Generated images Reconstructed
Figure 6: Ablation study on multi-view images generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
Generated images
K-Plane [7] ~2 hrs
No conf ~12 min
Full setting ~12 min
@@
We performed ablation studies to assess the influence of different components in our approach. We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [7]. The results corresponding to these settings can be found in Fig. 6 and Fig. 7.
@@
We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and fourth rows in Fig. 6. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation.
@@
As illustrated in the contrast between the third and fourth rows in Fig. 6, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
@@
The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise. This is evident in the contrast between the third and fourth rows in Fig. 7. The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality.
@@
Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [7]. In the comparison shown in Fig. 7, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes. Notably, our method surpasses K-Plane in speed, achieving a 10$\times$ improvement. Additionally, as emphasized in Sec. 4.3, our model enables real-time rendering with explicit point cloud geometry.
@@ 5 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model. Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 10 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.

### v2

@@ 4.3 Quantitative evaluation
Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. : Measured using code released by Consistent4D [13] with one A6000 GPU.
Consistent4D [13] Efficient4D(Ours)
- Comparison with ground truth on synthetic data
CLIP $\uparrow$ 0.87 0.92
LPIPS $\downarrow$ 0.16 0.14
- Comparison on all cases
CLIP $\uparrow$ 0.8471 0.8721
CLIP-T-f $\uparrow$ 0.9692 0.9702
CLIP-T-s $\uparrow$ 0.9658 0.9726
CLIP-T-b $\uparrow$ 0.9697 0.9707
Generation time $\downarrow$ 120 mins 6 mins
FPS $\uparrow$ ~20 ~1000
Explicit geometry
In Tab. 1, we compare Consistent4D and our method quantitatively. The metrics and data used are described in Sec. 4.1.
We draw several observations: (1) For synthetic data, our generation results produce better CLIP and LPIPS scores. (2) Our method achieves superior image quality and temporal smoothness on all cases. (3) Our method significantly accelerates the generation process, achieving over 20$\times$ speed improvement. (4) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization faster, and (2) each iteration in our method requiring much less time compared to Consistent4D. This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
@@
Input Ours Consistent4D [13]
Figure 6: Given only two input frames, our method is able to generate smooth dynamics. For each case, we show three internal images from two novel views.
@@ 4.4 Sparse input evaluation
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In such cases, we set $w=0.25$ in Eq. 2. As illustrated in Fig. 6, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency. In contrast, Consistent4D fails to operate effectively under such conditions.
@@ 4.5 Ablation studies
Input/Generated images Rendered Input/Generated images Rendered
Figure 7: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
K-Plane [8] PSNR: 33.50 Time: 2 hrs
Refer to caption Refer to caption
No conf PSNR: 32.13 Time: 6 mins
Refer to caption Refer to caption
Full setting PSNR: 32.64 Time: 6 mins
Refer to caption Refer to caption
@@
Table 2: Quantitative ablation study on all our modules mesured on 4 sequences: alpaca, astronaut, rabbit and spiderman.
Method CLIP $\uparrow$ CLIP-T-f $\uparrow$ CLIP-T-s $\uparrow$ CLIP-T-b $\uparrow$
Single-view 0.6684 - - -
No time-sync 0.8270 0.9258 0.9197 0.8819
No interp 0.8595 0.9409 0.9443 0.9336
No conf 0.8611 0.9658 0.9592 0.9611
Full setting 0.8702 0.9689 0.9673 0.9684
We performed ablation studies to assess the influence of different components in our approach. We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [8]. Note setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer. The results corresponding to these settings can be found in Fig. 7 and Fig. 8. Table 2 also gives quantitative evaluations.
@@
We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Fig. 7 and Tab. 2. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores.
@@
As illustrated in the contrast between the third and last rows in Fig. 7 and Tab. 2, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus delivering higher CLIP-T scores. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
@@
The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise. This is evident in the contrast between the “No conf” and “Full setting” results in Fig. 8 and Tab. 2. The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality and temporal smoothness.
@@
Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [8]. In the comparison shown in Fig. 8, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes. Notably, our method surpasses K-Plane in speed, achieving a 20$\times$ improvement. Additionally, as emphasized in Sec. 4.3, our model enables real-time rendering with explicit point cloud geometry.
@@ 5 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model. Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 20 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.